# Semi-parametric bulk and tail regression using spline-based neural networks

### v1

@@ 2.3 Assessing relative variable importance for SPQR
The SPQR framework quantifies the effect of the covariates on specific levels of the response distribution via the quantile function $Q_{\rm SPQR}(\tau|\mbox{\bf x})$, where $\tau\in(0,1)$ is the quantile level of interest. In previous work using SPQR (see e.g., Majumder et al., 2024b, ; Majumder and Reich, , 2023), covariate effects on $\tau$-quantiles have been measured using accumulative local effects (ALEs; Apley and Zhu, , 2020). We briefly outline ALEs and their link to relative variable importance quantification; we consider the effects of covariates on a generic differentiable function $g(\mbox{\bf x})$, where $\mbox{\bf x}=(x_{1},\ldots,x_{p})$ is the vector of covariates. The sensitivity of $g(\mbox{\bf x})$ to covariate $x_{j}$ is quantified by the partial derivative
@@
$\mbox{ALE}_{j}(x_{j};g)=\int_{z_{0,j}}^{x_{j}}\mathbb{E}[\dot{g}_{j}(x_{j})|x_{j}=z_{j}]\mathrm{d}z_{j},$
where $z_{0,j}$ is an approximate lower bound for $x_{j}$. In practice, $\mbox{ALE}_{j}(x_{j};g)$ is estimated by replacing the above integral with a partial sum, which can be used even when $g(\mbox{\bf x})$ is not differentiable. The $\mbox{ALE}_{j}(x_{j};g)$ scores provides the relative importance of a specific value of the $j$-th covariate, $x_{j}$, on the function $g(\cdot)$. Following Greenwell et al., (2018), we can measure the heterogeneity of the effect of covariate $X_{j}$ on $g(\cdot)$ by taking the standard deviation of $\mbox{ALE}_{j}(X_{j};g)$ with respect to the distribution of $X_{j}$. We thus define the variable importance (VI) scores for $X_{j}$ on $g(\cdot)$ by
$\mbox{VI}_{j}(g)$ $=\sqrt{\mbox{Var}_{X_{j}}[\mbox{ALE}_{j}(X_{j};g)]}.$
@@
3 Extremal semi-parametric quantile regression
@@ 3.1 Overview
In Section 3.2, we present the blended generalised Pareto distribution. In Section 3.3, we describe the extremal semi-parametric quantile regression (xSPQR) framework, which blends the SPQR representation for conditional densities with a deep GP regression model (as discussed in Section 1). Section 3.4 details inference for xSPQR, and Section 3.5 describes estimation of relative variable importance for the conditional upper-tails.
@@ 3.2 Blended generalised Pareto distribution
Castro-Camilo et al., (2022) introduced the blended generalised extreme value distribution (bGEV), which combines the lower (exponential) tail of a Gumbel distribution with the upper (heavy) tail of a Frechét distribution; Krakauer, (2024) extended the blended GEV to allow for bounded or light-tailed margins. The blended GEV has already been shown to be an effective model for deep extreme quantile regression; see, e.g., Richards and Huser, (2022) and Richards and Huser, (2024). We follow a similar approach to Castro-Camilo et al., (2022) and construct the blended GP distribution (bGP), which blends together a constituent “bulk” distribution, describing the body, with the GP distribution, for the upper-tails. We present here the bGP for an arbitrary bulk distribution, in the unconditional setting, i.e., we consider a continuous random variable $Y\geq 0$, but no covariate information. We discuss the conditional setting as a natural extension towards EVT-compliant density regression in Section 3.3, where we take the bulk distribution to be the SPQR representation of conditional densities, as defined in (3). This creates a conditional distribution function that exhibits great flexibility in the body, while retaining the exact GP upper-tails of standard parametric extreme quantile regression models.
Let $F(y|\mathcal{W})$, $f(y|\mathcal{W})$, and $Q(y|\mathcal{W})$ denote the distribution, density, and quantile functions, respectively, of a continuous random variable, which have estimable parameters contained in $\mathcal{W}$; here $F(y|\mathcal{W})$ and $f(y|\mathcal{W})$ are supported on the interval $[0,1]$ and referred to as the “bulk” distribution and density, respectively. We then let $F_{\rm GP}(y|u,\sigma_{u},\xi)$, $f_{\rm GP}(y|u,\sigma_{u},\xi)$, and $Q_{\rm GP}(y|u,\sigma_{u},\xi)$ denote the distribution, density, and quantile functions, respectively, of a generalised Pareto random variable with scale parameter $\sigma_{u}>0$, shape parameter $\xi\in\mathbb{R}$, and threshold $u\in\mathbb{R}$. The distribution function $F_{\rm GP}(y|u,\sigma_{u},\xi)$ is
@@
We define the blended GP distribution, denoted by bGP$(\mathcal{W},\xi),$ via its continuous distribution function $H(y|\mathcal{W},\xi)$, defined as
$H(y|\mathcal{W},\xi)=\begin{cases}F(y|\mathcal{W})^{1-p(y)}F_{\rm GP}(y|\tilde{u},\tilde{\sigma}_{u},\xi)^{p(y)},\quad&y>\tilde{u},\\ F(y|\mathcal{W}),\quad&y\leq\tilde{u},\\ \end{cases}$ (4)
where $p(y)\in[0,1]$ is a weighting function. The distribution function $H(y|\mathcal{W},\xi)$ inherits support from the constituent distribution functions; the lower bound of the support is zero, whilst the upper bound is infinite if $\xi\geq 0$ and $\tilde{u}-\tilde{\sigma}_{u}/\xi>0$, otherwise. Following Castro-Camilo et al., (2022), we take the weighting function to be
@@
$(\tilde{\sigma}_{u},\tilde{u})=\begin{cases}\left(\frac{\xi(a-b)}{(1-p_{a})^{-\xi}-(1-p_{b})^{-\xi}},a-\frac{(a-b)\{(1-p_{a})^{-\xi}-1\}}{(1-p_{a})^{-\xi}-(1-p_{b})^{-\xi}}\right),\quad\xi\neq 0,\\ \left(\frac{(a-b)}{\log(1-p_{b})-\log(1-p_{b})},a-\frac{(a-b)\{-\log(1-p_{a})\}}{\log(1-p_{b})-\log(1-p_{b})}\right),\quad\xi=0.\\ \end{cases}$ (5)
@@
The density function of a bGP$(\mathcal{W},\xi)$ random variable has the closed-form expression
@@
where $p^{\prime}(y)=(b-a)^{-1}f_{\rm Beta}\left(\frac{y-a}{b-a},c_{1},c_{2}\right)$ for $f_{\rm Beta}\left(\cdot,c_{1},c_{2}\right)$ the density of the Beta distribution with shape parameters $c_{1}$ and $c_{2}$. As discussed by Castro-Camilo et al., (2022), setting $c_{1},c_{2}>3$ ensures that the second derivative of the log-density is always continuous; we hereafter assume this to hold.
3.3 Tail density regression using xSPQR
@@
We now describe the extremal semi-parametric quantile regression framework, xSPQR. Remaining in the unconditional setting, we first create a flexible bGP distribution by constructing the constituent bulk distribution and density functions, $F(y|\mathcal{W})$ and $f(y|\mathcal{W})$, respectively, in (4), via the usual SPQR representation, i.e., we set $F(y|\mathcal{W})=F_{\rm SPQR}(y|\mathcal{W})$ (and $f(y|\mathcal{W})=f_{\rm SPQR}(y|\mathcal{W})$), which is constructed using a basis of $I$-splines (and $M$-splines) that are supported on the interval $[0,1]$, and with basis coefficients $\mathcal{W}=\{w_{k}\}^{K}_{k=1}$ (which here are not dependent on covariates). The original SPQR method took the basis to have equally-spaced knots on $[0,1]$ (Xu and Reich, , 2021). To reflect that xSPQR is designed for data with varying degrees of tail-heaviness (which may be covariate dependent), we instead place the knots at a sequence of quantiles with equally-spaced levels in $[0,1]$. The quantiles are estimated empirically using the training response data, and we ensure that knots are always placed at the boundaries of the unit interval.
@@
Figure 1 provides illustrative examples of the distribution and density functions for this flexible SPQR-based bGP model, alongside the constituent SPQR and GP distributions and density functions. We observe great flexibility in the bulk of the distribution, and the desired exact upper GP tails. As the hyper-parameters $p_{a}$ and $p_{b}$ have a clear interpretation as the beginning and end of the blending interval, we can dichotomize the density into the “bulk” for $y<a$, which is purely driven by SPQR, and the “upper-tail” for $y\geq b$, which is driven by both SPQR and $\xi$. The hyper-parameters $c_{1}$ and $c_{2}$ determine the relative weights placed on the constituent distribution functions in (4) during the blending interval. Comparing the first and third rows of Figure 1, we observe that a larger value of $c_{1}$, relative to $c_{2}$, places more weight on the SPQR distribution function when blending. Given that $F_{\rm SPQR}(\cdot|\mathcal{W})$ is substantially more flexible than $F_{\rm GP}(y|\tilde{u},\tilde{\sigma}_{u},\xi)$, we consider hereafter only the cases where $c_{1}\geq c_{2}$; for simplicity, we also set $c_{2}=5$ throughout.
@@
$H(y|\mathbf{x})=\begin{cases}F_{\rm SPQR}(y|\mathbf{x})^{1-p(y)}F_{\rm GP}(y|\tilde{u}(\mathbf{x}),\tilde{\sigma}_{u}(\mathbf{x}),\xi(\mathbf{x}))^{p(y)},\quad&y>\tilde{u}(\mathbf{x}),\\ F_{\rm SPQR}(y|\mathbf{x}),\quad&y\leq\tilde{u}(\mathbf{x}),\\ \end{cases}$ (7)
where $F_{\rm SPQR}(y|\mathbf{x})$ is as defined in (3). As in (5), the functions $\tilde{u}(\mathbf{x})$ and $\tilde{\sigma}_{u}(\mathbf{x})$ are deterministic functions of $p_{a}$, $p_{b}$, $a=Q_{\rm SPQR}(p_{a}|\mathbf{x})$, $b=Q_{\rm SPQR}(p_{b}|\mathbf{x}),$ and $\xi(\mathbf{x})$; note that the covariates $\mathbf{x}$ enter the conditional bGP distribution through both the constituent bulk distribution function, $F_{\rm SPQR}(y|\mathbf{x}),$ and the shape parameter of the GP distribution for the upper-tail; $\xi(\mbox{\bf x})$ can be represented as a flexible nonlinear function of x that is estimated alongside the weights of the conditional bulk distribution. We further note that the weighting function, $p(y)$, and its hyper-parameters, $p_{a}$, $p_{b}$, $c_{1}$, and $c_{2}$, are not dependent on covariates.
As outlined in (2.2), the basis weights $\mathcal{W}(\mathbf{x})=\{w_{k}(\mathbf{x})\}_{k=1}^{K}$ of the conditional density function in the SPQR framework, $f_{\rm SPQR}(y|\mathbf{x})$, are modelled as the final layer of a multi-layer perceptron; a softmax final activation layer is applied to ensure that the $K$-vector of weights must sum to one. For xSPQR, we must also model the shape parameter function, $\xi(\mathbf{x})$. A natural and parsimonious choice is to combine the modelling of $\xi(\mathbf{x})$ with that of $\mathcal{W}(\mathbf{x})$. To this end, we jointly model $(\xi(\mbox{\bf x}),\mathcal{W}(\mbox{\bf x}))$ as the final layer of a multi-layer perceptron (see Figure 2) with an appropriately chosen final activation layer; we replace the final layer of (2.2) by
$(\xi(\mbox{\bf x}),\mathcal{W}(\mbox{\bf x}))=\mbox{\bf x}^{(H+1)}=\mbox{softmax}^{*}(\mbox{\bf W}^{(H+1)}\mbox{\bf x}^{(H)}+\mbox{\bf b}^{(H+1)}),$
where the $\mbox{softmax}^{*}(\cdot)$ activation function applies a $\xi$-specific activation function to the first element of its input vector, and the usual $\mbox{softmax}(\cdot)$ activation to the last $K$ elements. The $\xi$-specific activation function should be chosen to ensure an appropriate range for $\xi(\mathbf{x})$, with examples including: exponential ($\xi(\mathbf{x})>0$), sigmoid ($\xi(\mathbf{x})\in(0,1)$), or tanh ($\xi(\mathbf{x})\in(-1,1)$), or scalings thereof; see discussion by Richards and Huser, (2024).
@@
Figure 2: Schematic of the underlying multi-layer percepton (MLP) of an xSPQR model with $H=2$ hidden layers. The input covariates are $\mathbf{x}=(x_{1},\dots,x_{p})$, and the output of the final $\mbox{softmax}^{*}(\cdot)$ layer is $(\xi(\mathbf{x}),w_{1}(\mathbf{x}),\dots,w_{K}(\mathbf{x})),$ where $w_{1}(\mathbf{x}),\dots,w_{K}(\mathbf{x})$ are the $K$ basis functions comprising the constituent SPQR density/distribution.
Hereon, we refer to this methodology of density regression as extremal SPQR, or xSPQR. The nomenclature emphasises the flexible distributional assumptions for our density regression model, but we note that alternative models for the body of the data could also be used in combination with the blended GP.

### v2

@@ 2.3 Assessing relative variable importance for SPQR
The SPQR framework quantifies the effect of the covariates on specific levels of the response distribution via the quantile function $Q_{\rm SPQR}(\tau|\mbox{\bf x})$, where $\tau\in(0,1)$ is the quantile level of interest. In previous work using SPQR (see e.g., Majumder et al., 2024b; Majumder and Reich, 2023), covariate effects on $\tau$-quantiles have been measured using accumulative local effects (ALEs; Apley and Zhu, 2020). We briefly outline ALEs and their link to relative variable importance quantification. We consider the effects of covariates on a generic differentiable function $g(\mbox{\bf x})$, where $\mbox{\bf x}=(x_{1},\ldots,x_{p})$ is the vector of covariates. The sensitivity of $g(\mbox{\bf x})$ to covariate $x_{j}$ is quantified by the partial derivative
@@
$\mbox{ALE}_{j}(x_{j};g)=\int_{z_{0,j}}^{x_{j}}\mathbb{E}\left[\dot{g}_{j}(x_{j})|x_{j}=z_{j}\right]\mathrm{d}z_{j},$
where $z_{0,j}$ is an approximate lower bound for $x_{j}$. In practice, $\mbox{ALE}_{j}(x_{j};g)$ is estimated by replacing the above integral with a partial sum, which can be used even when $g(\mbox{\bf x})$ is not differentiable. The $\mbox{ALE}_{j}(x_{j};g)$ scores provide the relative importance of a specific value of the $j$-th covariate, $x_{j}$, on the function $g(\cdot)$. Following Greenwell et al. (2018), we can measure the heterogeneity of the effect of covariate $X_{j}$ on $g(\cdot)$ by taking the standard deviation of $\mbox{ALE}_{j}(X_{j};g)$ with respect to the distribution of $X_{j}$. We thus define the variable importance (VI) score for $X_{j}$ on $g(\cdot)$ by
$\mbox{VI}_{j}(g)$ $=\sqrt{\mbox{Var}_{X_{j}}\left[\mbox{ALE}_{j}(X_{j};g)\right]}.$
@@
3 Semi-parametric quantile regression for extremes
@@ 3.1 Overview
In Section 3.2, we present the blended generalised Pareto distribution. In Section 3.3, we describe semi-parametric quantile regression for extremes (SPQRx), which blends the SPQR representation for conditional densities with a deep GP regression model (as discussed in Section 1). Section 3.4 details inference for SPQRx, and Section 3.5 describes estimation of relative variable importance for the conditional upper-tails.
@@ 3.2 Blended generalised Pareto distribution
Castro-Camilo et al. (2022) introduced the blended generalised extreme value distribution (bGEV), which combines the lower (exponential) tail of a Gumbel distribution with the upper (heavy) tail of a Frechét distribution; Krakauer (2024) extended the blended GEV to allow for bounded or light-tailed margins. The blended GEV has already been shown to be an effective model for deep extreme quantile regression; see, e.g., Richards and Huser (2026b) and Richards and Huser (2026a). We follow a similar approach to Castro-Camilo et al. (2022) and construct the blended GP distribution (bGP), which blends together a constituent “bulk” distribution, describing the body, with the GP distribution, for the upper-tails. We present here the bGP for an arbitrary bulk distribution, in the unconditional setting, i.e., we consider a continuous random variable $Y\geq 0$, but no covariate information. We discuss the conditional setting as a natural extension towards EVT-compliant density regression in Section 3.3, where we take the bulk distribution to be the SPQR representation of conditional densities, as defined in (3). This creates a conditional distribution function that exhibits great flexibility in the body, while retaining the exact GP upper-tails of standard parametric extreme quantile regression models.
Let $F(y|\mathcal{W})$, $f(y|\mathcal{W})$, and $Q(y|\mathcal{W})$ denote the distribution, density, and quantile functions, respectively, of a continuous random variable, which have estimable parameters contained in $\mathcal{W}$. Here $F(y|\mathcal{W})$ and $f(y|\mathcal{W})$ are supported on the interval $[0,1]$ and referred to as the “bulk” distribution and density, respectively. We then let $F_{\rm GP}(y|u,\sigma_{u},\xi)$, $f_{\rm GP}(y|u,\sigma_{u},\xi)$, and $Q_{\rm GP}(y|u,\sigma_{u},\xi)$ denote the distribution, density, and quantile functions, respectively, of a GP random variable with scale parameter $\sigma_{u}>0$, shape parameter $\xi\in\mathbb{R}$, and threshold $u\in\mathbb{R}$. The distribution function $F_{\rm GP}(y|u,\sigma_{u},\xi)$ is
@@
[…]
where $p(y)\in[0,1]$ is a weighting function. The distribution function $H(y|\mathcal{W},\xi)$ inherits support from the constituent distribution functions; the lower bound of the support is zero, whilst the upper bound is infinite if $\xi\geq 0$ and $\tilde{u}-\tilde{\sigma}_{u}/\xi>0$, otherwise. Following Castro-Camilo et al. (2022), we take the weighting function to be
@@
where $F_{\rm Beta}\left(\cdot,c_{1},c_{2}\right)$ denotes the distribution function of a Beta($c_{1},c_{2})$ random variable with shape parameters $c_{1}>0$ and $c_{2}>0$. We blend the density functions $F(\cdot|\mathcal{W})$ and $F_{\rm GP}(\cdot|u,\sigma_{u},\xi)$ within the interval $[a,b]$, where the bounds are taken to be the $p_{a}$ and $p_{b}$ quantiles of $F(\cdot)$. That is, for $p_{a},p_{b}\in(0,1)$ with $p_{a}<p_{b}$, we take $a:=Q(p_{a}|\mathcal{W})$ and $b:=Q(p_{b}|\mathcal{W})$. To ensure continuity of $H(\cdot|\mathcal{W},\xi)$, we require $F(a|\mathcal{W})=F_{\rm GP}(a|\tilde{u},\tilde{\sigma}_{u},\xi)$ and $F(b|\mathcal{W})=F_{\rm GP}(b|\tilde{u},\tilde{\sigma}_{u},\xi)$. This is achieved by setting
@@
[…]
@@
where $p^{\prime}(y)=(b-a)^{-1}f_{\rm Beta}\left(\frac{y-a}{b-a},c_{1},c_{2}\right)$ for $f_{\rm Beta}\left(\cdot,c_{1},c_{2}\right)$ the density of the Beta distribution with shape parameters $c_{1}$ and $c_{2}$. As discussed by Castro-Camilo et al. (2022), setting $c_{1},c_{2}>3$ ensures that the second derivative of the log-density is always continuous; we hereafter assume this to hold.
While $h(y|\mathcal{W},\xi)$ in (6) is guaranteed to be continuous and to integrate to one, it is not guaranteed to be positive for all $y$, i.e., not a valid density function. There exist pathological choices of $p(\cdot)$ such that, given fixed $f(\cdot|\mathcal{W})$ and $\xi$, $h(y|\mathcal{W},\xi)<0$ for $y\in[a,b]$} where $[a,b]$ is the blending interval. We note that this occurs only if $F(\cdot|\mathcal{W})$ is right-skewed and $F_{\rm GP}(\cdot|\xi)$ is relatively lighter-tailed, and $F(y|\mathcal{W})>F_{\rm GP}(y|\mathcal{W})$ for some $y\in[a,b]$}. In practice, this is very unlikely to be an issue as $f(\cdot|\mathcal{W})$ and $\xi$ are estimated jointly, and we consider this to be only a theoretical limitation of the blended GP formulation. We further note that this is fully precluded by considering $c_{2}/c_{1}\rightarrow\infty$, which corresponds to the case where $H(y|\mathcal{W},\xi)=F(y|\mathcal{W})$ for $y\in[a,b]$ (or similarly for $c_{1}/c_{2}\rightarrow\infty$ and $H(y|\mathcal{W},\xi)=F_{\rm GP}(y|\xi)$). Although we found no evidence of this problem occurring during inference of any models discussed hereafter, we do provide further discussion of this problem, and mitigating its occurrence during inference, in Appendix B.
3.3 Tail density regression using SPQRx
@@
We now describe semi-parametric quantile regression for extremes, SPQRx. Remaining in the unconditional setting, we first create a flexible bGP distribution by constructing the constituent bulk distribution and density functions, $F(y|\mathcal{W})$ and $f(y|\mathcal{W})$, respectively, in (4), via the usual SPQR representation, i.e., we set $F(y|\mathcal{W})=F_{\rm SPQR}(y|\mathcal{W})$ (and $f(y|\mathcal{W})=f_{\rm SPQR}(y|\mathcal{W})$), which is constructed using a basis of $I$-splines (and $M$-splines) that are supported on the interval $[0,1]$, and with basis coefficients $\mathcal{W}=\{w_{k}\}^{K}_{k=1}$ (which here are not dependent on covariates). The original SPQR method took the basis to have equally-spaced knots on $[0,1]$ (Xu and Reich, 2021). To reflect that SPQRx is designed for data with varying degrees of tail-heaviness (which may be covariate dependent), we instead place the knots at a sequence of quantiles with equally-spaced levels in $[0,1]$. The quantiles are estimated empirically using the training response data, and we ensure that knots are always placed at the boundaries of the unit interval.
@@
Figure 1 provides illustrative examples of the distribution and density functions for this flexible SPQR-based bGP model, alongside the constituent SPQR and GP distributions and density functions. We observe great flexibility in the bulk of the distribution, and the desired exact upper GP tails. As the hyper-parameters $p_{a}$ and $p_{b}$ have a clear interpretation as the beginning and end of the blending interval, we can dichotomize the density into the “bulk” for $y<a$, which is purely driven by SPQR, and the “upper-tail” for $y\geq b$, which is driven by both SPQR and $\xi$. The hyper-parameters $c_{1}$ and $c_{2}$ determine the relative weights placed on the constituent distribution functions in (4) during the blending interval. Comparing the first and third rows of Figure 1, we observe that a larger value of $c_{1}$, relative to $c_{2}$, places more weight on the SPQR distribution function when blending. Given that $F_{\rm SPQR}(\cdot|\mathcal{W})$ is substantially more flexible than $F_{\rm GP}(y|\tilde{u},\tilde{\sigma}_{u},\xi)$, we consider hereafter only the cases where $c_{1}\geq c_{2}$. For simplicity, we also set $c_{2}=5$ throughout.
Figure 1: Illustrative examples of the blended GP distribution (left) and density (right) functions, with the constituent distribution modelled via SPQR. For a fixed random seed, we generate $K$ basis coefficients and construct the SPQR density and distribution functions (green curves). We then find $a=Q_{\rm SPQR}(p_{a}|\mathcal{W})$ and $b=Q_{\rm SPQR}(p_{b}|\mathcal{W})$, where $p_{a}=0.25$ and $p_{b}=0.9$. The values of $a$ and $b$ are denoted by the orange horizontal lines. Then, we find the required GP distribution function (red) to satisfy continuity of the bGP($\mathcal{W},\xi)$ distribution function; its resulting distribution and density functions are provided in black. The weighting function shape parameter $c_{2}$ is fixed to $c_{2}=5$ across all panels.
@@
where $F_{\rm SPQR}(y|\mathbf{x})$ is as defined in (3). As in (5), the functions $\tilde{u}(\mathbf{x})$ and $\tilde{\sigma}_{u}(\mathbf{x})$ are deterministic functions of $p_{a}$, $p_{b}$, $a=Q_{\rm SPQR}(p_{a}|\mathbf{x})$, $b=Q_{\rm SPQR}(p_{b}|\mathbf{x}),$ and $\xi(\mathbf{x})$. Note that the covariates $\mathbf{x}$ enter the conditional bGP distribution through both the constituent bulk distribution function, $F_{\rm SPQR}(y|\mathbf{x}),$ and the shape parameter $\xi(\mbox{\bf x})$ of the GP distribution for the upper-tail. The latter is estimated alongside the weights of the conditional bulk distribution. We further note that the weighting function, $p(y)$, and its hyper-parameters, $p_{a}$, $p_{b}$, $c_{1}$, and $c_{2}$, are not dependent on covariates.
The SPQRx conditional quantile function, $Q_{\rm SPQRx}(\tau|\mathbf{x}):=H^{-1}(\tau|\mathbf{x})$, can be evaluated for any probability level $\tau\in[0,1]$. Although it does not have a closed form expression, its numerical evaluation is fast: for $\tau\leq p_{a}$, $Q_{\rm SPQRx}(\tau|\mathbf{x})=Q_{\rm SPQR}(\tau|\mathbf{x})$, which can be quickly approximated using interpolation; for $\tau\geq p_{b}$, $Q_{\rm SPQRx}(\tau|\mathbf{x})=Q_{\rm GP}(\tau|\tilde{u}(\mathbf{x}),\tilde{\sigma}_{u}(\mathbf{x}),\xi(\mathbf{x}))$ has a closed form expression; for $\tau\in[p_{a},p_{b}]$, $Q_{\rm SPQRx}(\tau|\mathbf{x})$ must be solved numerically but the region $p_{a},p_{b}]$ is typically small. The continuous semi-parametric form of the conditional quantile function allows for joint estimation of quantiles at a sequence of levels $\tau$. Thus, as long as $Q_{\rm SPQRx}$ is monotonically increasing, our method does not suffer from the crossing problems of standard quantile regression approaches (see, e.g., Saleh and Saleh, 2021).
As outlined in (2.2), the basis weights $\mathcal{W}(\mathbf{x})=\{w_{k}(\mathbf{x})\}_{k=1}^{K}$ of the conditional density function in the SPQR framework, $f_{\rm SPQR}(y|\mathbf{x})$, are modelled as the final layer of a multi-layer perceptron. A softmax final activation layer is applied to ensure that the $K$-vector of weights must sum to one. For SPQRx, we must also model the shape parameter function, $\xi(\mathbf{x})$. A natural and parsimonious choice is to combine the modelling of $\xi(\mathbf{x})$ with that of $\mathcal{W}(\mathbf{x})$. To this end, we jointly model $(\xi(\mbox{\bf x}),\mathcal{W}(\mbox{\bf x}))$ as the final layer of a multi-layer perceptron (see Figure 2) with an appropriately chosen final activation layer. We replace the final layer of (2.2) by
$\left(\xi(\mbox{\bf x}),\mathcal{W}(\mbox{\bf x})\right)=\mbox{\bf x}^{(H+1)}=\mbox{softmax}^{*}\left(\mbox{\bf W}^{(H+1)}\mbox{\bf x}^{(H)}+\mbox{\bf b}^{(H+1)}\right),$
where the $\mbox{softmax}^{*}(\cdot)$ activation function applies a $\xi$-specific activation function to the first element of its input vector, and the usual $\mbox{softmax}(\cdot)$ activation to the last $K$ elements. The $\xi$-specific activation function should be chosen to ensure an appropriate range for $\xi(\mathbf{x})$, with examples including: exponential ($\xi(\mathbf{x})>0$), sigmoid ($\xi(\mathbf{x})\in(0,1)$), or tanh ($\xi(\mathbf{x})\in(-1,1)$); see discussion by Richards and Huser (2026a).
@@
Figure 2: Schematic of the underlying multi-layer percepton (MLP) of an SPQRx model with $H=2$ hidden layers. The input covariates are $\mathbf{x}=(x_{1},\dots,x_{p})$, and the output of the final $\mbox{softmax}^{*}(\cdot)$ layer is $(\xi(\mathbf{x}),w_{1}(\mathbf{x}),\dots,w_{K}(\mathbf{x})),$ where $w_{1}(\mathbf{x}),\dots,w_{K}(\mathbf{x})$ are the $K$ basis functions comprising the constituent SPQR density/distribution.
Throughout, we refer to this methodology of density regression as SPQR for extremes, or SPQRx. The nomenclature emphasises the flexible distributional assumptions for our density regression model, but we note that alternative models for the body of the data could also be used in combination with the blended GP.